# Online SLA Decomposition: Enabling Real-Time Adaptation to Evolving Systems

### v2

@@ Abstract
When a network slice spans multiple domains, each domain must uphold the End-to-End (E2E) Service Level Agreement (SLA) associated with the slice. This requires decomposing the E2E SLA into partial SLAs for each domain. In a two-level network slicing management system with an E2E orchestrator and local controllers, we propose an online learning-decomposition framework that dynamically updates risk models using recent feedback. This approach utilizes online gradient descent and FIFO memory buffers to enhance stability and robustness. Our empirical study shows the proposed framework outperforms state-of-the-art static methods, offering more accurate and resilient SLA decomposition under varying conditions and sparse data.
@@ I Introduction
The fifth generation (5G) of mobile communication technology introduced a versatile, multi-service network designed to support a wide range of vertical industries with a diverse set of service requirements. In 5G and beyond, network slicing plays a pivotal role by enabling the establishment and management of multiple End-to-End (E2E) logical networks. These slices are built on shared infrastructure and are specifically customized to meet the particular requirements of a given service, which is outlined in Service Level Agreements (SLAs). SLAs function as contracts between service providers and tenants, defining the expected Quality of Service (QoS) through well-defined, measurable benchmarks known as Service-Level Objectives (SLOs). These objectives encompass various performance metrics such as data throughput, latency, reliability, and security, among others. A single network slice may traverse multiple segments of the network, including access, core, and transport networks, and it may involve collaboration between different operators and infrastructure providers. To ensure that the service meets the agreed-upon SLOs across these domains, it is essential to adjust the service parameters accordingly. As a result, the E2E SLA linked to a network slice must be broken down into specific SLOs for each domain, allowing for effective resource allocation within each segment. The authors in [1] highlight the importance of E2E SLA decomposition for resource allocation, while [2] indicates that AI-assisted SLA decomposition is key to automating complex 6G business. Following the scenarios in [3, 4], in this paper we consider a two-level management architecture with an E2E service orchestrator handling network service lifecycle management, and local domain controllers managing slice instantiation within their domains (Fig. 1). The orchestrator determines the SLA decomposition for incoming service requests, while domain controllers handle admission control and resource allocation. We assume the orchestrator lacks real-time knowledge of the infrastructure state but has access to historical feedback (i.e., request acceptance or rejection) from each domain. This allows the orchestrator to make informed decisions using domain-specific risk models based on the available data.
@@
Several SLA decomposition methods employing heuristics have been studied [1]. Authors in [5] present an E2E SLA decomposition system that applies supervised machine learning to break down E2E SLAs into access, transport, and core SLOs. In our previous work [3, 4], we tackle the problem with a two-step approach, which is a combination of machine learning and optimization-based solution.
@@
    We conduct an ablation study to examine the contribution and importance of individual component of the proposed framework, and compare with the state-of-the-art offline learning approach.
The rest of this paper is organized as follows; Section II presents background information and problem formulation on the SLA decomposition. In Section 2 we present the proposed online learning-decomposition framework. Section IV details the experimental setup, followed by Section V, where we assess the performance of the proposed framework. Finally, Section VI presents our conclusions.
@@ II-A SLA Decomposition and Risk Models
An E2E SLA, denoted by $s_{e2e}$, is a collection of Service Level Objectives (SLOs) linked to specific performance indicators. The SLO vector outlines the SLA’s performance requirements in a sequential manner. For example, an SLA encompassing E2E delay and throughput is expressed as $s_{e2e}=(\tau_{e2e},\theta_{e2e})$. This implies that the network slice must operate in a way that meets constraints imposed by $\tau_{e2e}$ for delay and $\theta_{e2e}$ for throughput, ensuring $\tau\leq\tau_{e2e}$, and $\theta\geq\theta_{e2e}$. Considering a network slice distributed across $N$ domains (where n ranges from 1 to $N$), we introduce $s_{n}$ to represent the SLOs of the $n$-th domain. The relationship between individual domain SLOs and the overall E2E objective $s_{e2e}$ is defined by $s_{e2e}=G(s_{1},s_{2},...,s_{N})$. For instance, the E2E delay is the sum of all delays for the involved domains, while the E2E throughput is determined by the lowest throughput across all domains. Mathematically, this is represented as $\tau_{e2e}=\Sigma_{n=1}^{N}\tau_{n}$ and $\theta_{e2e}=\min\{\theta_{1},\theta_{2},...,\theta_{N}\}$.
@@ II-C Dynamicity
The dynamicity of the system is crucial for accurately managing SLAs in network domains. Multiple factors contribute to this dynamicity, including traffic intensity fluctuations, changes in user behavior, varying network conditions and security threats. As these factors change, the acceptance probability of requests within a domain varies, impacting overall network performance. Particularly, the acceptance probability becomes time-dependent. At a discrete time step $t$, the acceptance probability (in a single domain) of the SLA request $s$ is denoted as $P_{t}(s)$, which is influenced by the state $\omega_{t}$ of the domain, as stated in Section II-B. Given that $\omega_{t}$ evolves over time, the acceptance probability $P_{t}(s)$ also varies, even for the same SLOs. This behavior necessitates continuous learning and adaptation of the risk models. Dynamic systems require periodic updates to risk models based on the recent feedback to ensure these models remain relevant and accurate. In the next section, we introduce a novel framework that allows the system to adapt to varying conditions and maintain near-optimal performance.
@@ III Approaches
In this section, we present the proposed framework Real-time Adaptative DEcomposition (RADE).
@@ III-B Online Learning Framework
To capture the dynamic nature of the environment, where the decision-making process of controllers evolves over time, it is essential to constantly update risk models based on the recent feedback. To this end, we propose an online learning-decomposition framework RADE, which is capable of running stable update as well as providing resilience against noisy samples. The following are the key components of RADE: Base model. Following the design in [4], the base model is an NN. To account for monotonicity, we specifically choose Absolute Weight Transformation (AWET) model that shows prominent performance while requiring the least training effort, as suggested in [4].
@@
By updating the risk models online, we ensure that the decomposition remains responsive and resilient to changes. This continuous adaptation process not only enhances the accuracy of the model but also improves the overall performance and reliability of the system. Alg. 1 details the steps of the online update and decomposition mechanism. In the subsequent sections, we will present the simulation environment for evaluations, followed by empirical results demonstrating its effectiveness.
@@ V Results
Fig. 3 presents the average E2E acceptance probability across different arrival rates for four methods: Static, RADE*, RADE, and OPT. The arrival rates vary between 0.3, 0.5, and 0.7, representing different traffic intensities in the system. Across all arrival rates, the Static method consistently exhibits the lowest average E2E acceptance probability. This suggests that the Static method’s one-off trained risk models are less adaptable to varying traffic conditions, leading to sub-optimal performance, particularly under higher traffic loads. RADE* shows improved performance over the Static method, which indicates that dynamically updating the risk models, even without the FIFO memory buffer, leads to a better adaptability than the Static approach. The RADE method, which includes the FIFO memory buffer for maintaining historical feedback, outperforms both Static and RADE*, highlighting the importance of the use of the FIFO buffer in enhancing the robustness and stability of the framework, particularly under varying traffic conditions. It is interesting to see that the performance of RADE* and RADE becomes nearly identical at higher arrival rates because, under heavy traffic, the system receives a large volume of recent feedback. This plenty of fresh data diminishes the impact of the FIFO memory buffer in RADE, as the most recent observations dominate the learning process. As expected, the OPT method achieves the highest performance, which serves as a theoretical upper bound. In contrast, the Random method performs significantly worse than the other approaches in general, with results of 0.42, 0.38, and 0.35 for arrival rates of 0.3, 0.5, and 0.7, respectively. Due to the consistently low performance of the Random method, we exclude its results from all subsequent figures to focus on the more meaningful comparisons.
@@
This paper has introduced an online learning-decomposition framework for SLA management in dynamic multi-domain environments. The proposed framework updates domain-specific risk models in real-time based on recent feedback, hence effectively addressing the challenges posed by the dynamicity of real-world systems. With the use of online update and FIFO memory buffers, the proposed framework enhances the stability and robustness of SLA decomposition, ensuring that the system is able to adapt promptly to changes in network conditions. Empirical results demonstrate that our framework outperforms state-of-the-art static methods, offering a more resilient and accurate decomposition of SLAs, even with limited data availability. In future work, we will investigate the potential of Deep Reinforcement Learning (DRL) to further enhance the SLA decomposition process. DRL could be a promising choice for this task due to its ability to learn optimal long-term policies through interaction with dynamic environments, which makes it an attractive option for improving SLA management in next-generation networks.

### v3

@@ Abstract
When a network slice spans multiple technology domains, it is crucial for each domain to uphold the End-to-End (E2E) Service Level Agreement (SLA) associated with the slice. Consequently, the E2E SLA must be properly decomposed into partial SLAs that are assigned to each domain involved. In a network slice management system with a two-level architecture, comprising an E2E service orchestrator and local domain controllers, we consider that the orchestrator has access solely to historical data regarding the responses of local controllers to previous requests, and this information is used to construct a risk model for each domain. In this study, we extend our previous work by investigating the dynamic nature of real-world systems and introducing an online learning-decomposition framework to tackle the dynamicity. We propose a framework that periodically updates the risk models based on the most recent feedback. This approach leverages key components such as online gradient descent and FIFO memory buffers, which enhance the stability and robustness of the overall process. Our empirical study on an analytic model-based simulator demonstrates that the proposed framework outperforms the state-of-the-art static approach, providing more accurate and resilient SLA decomposition even under varying conditions and limited data scenarios.
@@ I Introduction
The fifth generation (5G) of mobile communication technology introduced a versatile, multi-service network designed to support a wide range of vertical industries with a diverse set of service requirements. In 5G and beyond, network slicing plays a pivotal role by enabling the establishment and management of multiple End-to-End (E2E) logical networks. These slices are built on shared infrastructure and are specifically customized to meet the particular requirements of a given service, which is outlined in Service Level Agreements (SLAs). SLAs function as contracts between service providers and tenants, defining the expected Quality of Service (QoS) through well-defined, measurable benchmarks known as Service-Level Objectives (SLOs). These objectives encompass various performance metrics such as data throughput, latency, reliability, and security, among others. A single network slice may traverse multiple segments of the network, including (radio) access, transport, and core networks, and it may involve collaboration between different operators and infrastructure providers. To ensure that the service meets the agreed-upon SLOs across these domains, it is essential to adjust the service parameters accordingly. As a result, the E2E SLA linked to a network slice must be broken down into specific SLOs for each domain, allowing for effective resource allocation within each segment. The problem has been described in [1, 2, 3]. In [1], authors suggested decomposing the end-to-end requirement to the transport network requirement is one of the key issues in network slice requirement mapping. Authors in [2] outlined how slices are managed through lifecycle automation, orchestration, and real-time monitoring to ensure SLA compliance across different domains. In [3], authors elaborate that E2E QoS depends on SLA parameters and that transport resources must align with the heterogeneous QoS needs of different slices. The authors in [4] highlighted the importance of E2E SLA decomposition for resource allocation, while [5] indicates that AI-assisted SLA decomposition is key to automating complex 6G business. Following the scenarios in [6, 7], in this paper we consider a two-level management architecture with an E2E service orchestrator handling network service lifecycle management, and local domain controllers managing slice instantiation within their domains, as shown in Fig. 1. The orchestrator determines the SLA decomposition for incoming service requests, while domain controllers handle admission control and resource allocation. We assume the orchestrator lacks real-time knowledge of the infrastructure state but has access to historical feedback (i.e., request acceptance or rejection) from each domain. This allows the orchestrator to make informed decisions using domain-specific risk models based on the available data.
@@
Some studies [8, 9, 10] have proposed prediction-based approaches for SLAs management, although they do not explicitly address E2E SLAs decomposition problem. In [8], authors proposed the use of a mapping layer, which supervises the network over a service area and manages the allocation of radio resources to slices to guarantee their target service requirements. Authors in [9] proposed a SLA-constrained optimization using Deep Learning (DL) to estimate the required resources based on the traffic per slice. Another work [10] employs a context approach using graph representations for SLA violations prediction in Cloud Computing. Moreover, several SLA decomposition methods employing heuristics have been studied [4]. Authors in [11] present an E2E SLA decomposition system that applies supervised machine learning to break down E2E SLAs into access, transport, and core SLOs. In our previous work [6, 7], we tackle the problem with a two-step approach, which is a combination of machine learning and optimization-based solution.
@@
    We conduct an ablation study to examine the contribution and importance of individual component of the proposed framework, and compare with the state-of-the-art offline learning approach, followed by a complexity analysis.
The rest of this paper is organized as follows; Section II presents background information and problem formulation on the SLA decomposition. In Section 2 we present the proposed online learning-decomposition framework. Section IV details the experimental setup, followed by Section V, where we assess the performance and model complexity of the proposed framework. Finally, Section VI presents our conclusions.
@@ II-A SLA Decomposition and Risk Models
An E2E SLA, denoted by $s_{e2e}$, is a collection of SLOs linked to specific performance indicators. The SLO vector outlines the SLA’s performance requirements in a sequential manner. For example, an SLA encompassing E2E delay and throughput is expressed as $s_{e2e}=(\tau_{e2e},\theta_{e2e})$. This implies that the network slice must operate in a way that meets constraints imposed by $\tau_{e2e}$ for delay and $\theta_{e2e}$ for throughput, ensuring $\tau\leq\tau_{e2e}$, and $\theta\geq\theta_{e2e}$. Considering a network slice distributed across $N$ domains (where n ranges from 1 to $N$), we introduce $s_{n}$ to represent the SLOs of the $n$-th domain. The relationship between individual domain SLOs and the overall E2E objective $s_{e2e}$ is defined by $s_{e2e}=G(s_{1},s_{2},...,s_{N})$. For instance, the E2E delay is the sum of all delays for the involved domains, while the E2E throughput is determined by the lowest throughput across all domains. Mathematically, this is represented as $\tau_{e2e}=\Sigma_{n=1}^{N}\tau_{n}$ and $\theta_{e2e}=\min\{\theta_{1},\theta_{2},...,\theta_{N}\}$.
@@ II-C Dynamicity
The dynamicity of the system is crucial for accurately managing SLAs in network domains. Multiple factors contribute to this dynamicity, including traffic intensity fluctuations, changes in user behavior, varying network conditions and security threats. As these factors change, the acceptance probability of requests within a domain varies, impacting overall network performance. Particularly, the acceptance probability becomes time-dependent. At a discrete time step $t$, the acceptance probability (in a single domain) of the SLO request $s$ is denoted as $P_{t}(s)$, which is influenced by the state $\omega_{t}$ of the domain, as stated in Section II-B. Given that $\omega_{t}$ evolves over time, the acceptance probability $P_{t}(s)$ also varies, even for the same SLOs. This behavior necessitates continuous learning and adaptation of the risk models. Dynamic systems require periodic updates to risk models based on the recent feedback to ensure these models remain relevant and accurate. In the next section, we introduce a novel framework that allows the system to adapt to varying conditions and maintain near-optimal performance.
@@ III Approaches
In this section, we present the proposed framework Real-time Adaptive DEcomposition (RADE).
@@ III-B Online Learning Framework
To capture the dynamic nature of the environment, where the decision-making process of controllers evolves over time, it is essential to constantly update risk models based on the recent feedback. To this end, we propose an online learning-decomposition framework RADE, which is capable of running stable update as well as providing resilience against noisy samples. The following are the key components of RADE: Base model. Following the design in [7], the base model is an NN. To account for monotonicity described in (4), we employ Absolute Weight Transformation (AWET) approach that shows prominent performance. AWET ensures that the weights remain non-negative, a sufficient condition for an NN to be monotonic, while still allowing the model’s weights to be optimized freely during training.
@@
Alg. 1 details the steps of the online update and decomposition mechanism. The RADE framework operates in a time-stepped manner, where two key processes—learning and inference—run in parallel during each time step. At the beginning of each step, the framework receives feedback from the network regarding the performance of recent service requests. This feedback, represented as a set $\mathcal{K}_{t}$, is pushed into a FIFO memory buffer $\mathcal{M}$ to ensure that only the most recent and relevant data are stored. This buffer helps balance between old and new feedback. Using this stored feedback, the framework updates its risk models $\mathcal{F}_{\theta}$ through an online gradient descent process, adjusting the parameters based on the gradient of the loss function defined in (3). These continuously updated models help the system stay adaptive to changing network conditions. Simultaneously, the framework processes a set of new service requests $\mathcal{R}_{t}$, each accompanied by an E2E SLA $\mathcal{S}_{e2e}$ that defines the expected performance. The E2E SLAs are then decomposed into partial SLAs $\mathcal{S}_{partial}$ for individual network domains by leveraging the latest risk models and optimizing (1) and (2). The decomposition process ensures that the E2E SLA are allocated across the network domains in an optimal manner such that the overall E2E acceptance probability is maximized. In the subsequent sections, we will present the simulation environment for evaluations, followed by empirical results.
@@ V Results
V-A Performance
Fig. 3 presents the average E2E acceptance probability across different arrival rates for four methods: Static, RADE*, RADE, and OPT. The arrival rates vary between 0.3, 0.5, and 0.7, representing different traffic intensities in the system. Across all arrival rates, the Static method consistently exhibits the lowest average E2E acceptance probability. This suggests that the Static method’s one-off trained risk models are less adaptable to varying traffic conditions, leading to sub-optimal performance, particularly under higher traffic loads. RADE* shows improved performance over the Static method, which indicates that dynamically updating the risk models, even without the FIFO memory buffer, leads to a better adaptability than the Static approach. The RADE method, which includes the FIFO memory buffer for maintaining historical feedback, outperforms both Static and RADE*, highlighting the importance of the use of the FIFO buffer in enhancing the robustness and stability of the framework, particularly under low arrival rate conditions. It is interesting to see that the performance of RADE* and RADE becomes nearly identical at higher arrival rates because, under heavy traffic, the system receives a large volume of recent feedback. This plenty of fresh data diminishes the impact of the FIFO memory buffer in RADE, as the most recent observations dominate the learning process. As expected, the OPT method achieves the highest performance, which serves as a theoretical upper bound. In contrast, the Random method performs significantly worse than the other approaches in general, with results of 0.42, 0.38, and 0.35 for arrival rates of 0.3, 0.5, and 0.7, respectively. Due to the consistently low performance of the Random method, we exclude its results from all subsequent figures to focus on the more meaningful comparisons.
@@ V-A Performance
V-B Complexity analysis
The comparison of time complexity is provided in Table I. Three methods (Static, RADE* and RADE) are analyzed regarding both the complexity of base model training and the decomposition processes, which run in parallel. Assume the architecture shared by all methods has a training complexity $O(f(D))$ for a single iteration, and $O(g(D))$ for decomposition, where $D$ is the size of the input data, and $f$, $g$ are functions dependent on the architecture. For the Static method, the time complexity is simply $O(g(D))$, since there is no additional overhead incurred by updates. For RADE*, the total complexity for $N$-iteration training is $O(Nf(D))$, and the overall complexity is $O(\max(Nf(D),g(D)))$. Similarly, RADE behaves like RADE* but with the additional use of the FIFO memory buffer. Assume the FIFO memory buffer can hold up to $Q$ data points, the overall complexity is $O(\max(Nf(Q),g(D)))$. Typically, $Q$ is larger than $D$ to ensure effective utilization of the FIFO memory buffer. In summary, the Static method is the most efficient in terms of time complexity, followed by RADE*. RADE, however, is the most computationally intensive method, primarily due to the additional overhead of performing online training on data stored in the FIFO memory buffer.
Method Time complexity
Static $O(g(D))$
RADE* $O(\max(Nf(D),g(D)))$
RADE $O(\max(Nf(Q),g(D)))$
TABLE I: Comparison of time complexity.
@@ V-B Complexity analysis
This paper has introduced an online learning-decomposition framework for SLA management in dynamic multi-domain environments. The proposed framework updates domain-specific risk models in real-time based on recent feedback, hence effectively addressing the challenges posed by the dynamicity of real-world systems. With the use of online update and FIFO memory buffers, the proposed framework enhances the stability and robustness of SLA decomposition, ensuring that the system is able to adapt promptly to changes in network conditions. Empirical results demonstrate that our framework outperforms state-of-the-art static methods, offering a more resilient and accurate decomposition of SLAs, even with limited data availability. In future work, we will explore the potential of Deep Reinforcement Learning (DRL) to further refine the SLA decomposition process. DRL is particularly promising for this task due to its capacity to optimize long-term rewards, which is crucial in dynamic environments where immediate feedback may not fully capture the impact of decisions. By learning to cope with delayed rewards, DRL can develop strategies that anticipate the cumulative effects of actions over time, making it an ideal approach for managing SLA decomposition in complex, evolving network conditions.